# On the existence of forward self-similar solutions to the two-dimensional Navier-Stokes equations

### v1

@@ 3.3. Weighted $H^{2}$-Estimates for $\nabla v_{\textnormal{re}}$
$\begin{cases}-\Delta V+\overline{V}\cdot\nabla V-\dfrac{1}{2}(y\cdot\nabla V+V)+\nabla P=f&\text{in }\mathbb{R}^{2},\\ \operatorname{div}V=0.\end{cases}$
@@
(3.11) $-\Delta v_{\text{re}}$ $=\frac{1}{2}y\cdot\nabla v_{\text{re}}+\frac{1}{2}v_{\text{re}}-\mathbb{P}(v_{\text{re}}\cdot\nabla v_{\text{re}}-v_{0}\cdot\nabla v_{\text{re}}-v_{\text{re}}\cdot\nabla v_{0}-v_{0}\cdot\nabla v_{0})$
@@
Let $V_{k}=\sqrt{1+|y|^{2}}\partial_{y_{k}}v_{\text{re}}$ and $q_{k}=\sqrt{1+|y|^{2}}\partial_{y_{k}}q$, we can find that
@@
$\quad+v_{\text{re}}\cdot\frac{y}{\sqrt{1+|y|^{2}}}\partial_{y_{k}}v_{\text{re}}-\frac{1}{2}V_{k}-\frac{|y|^{2}}{2\sqrt{1+|y|^{2}}}\partial_{y_{k}}v_{\text{re}}$
$\quad-2\frac{y}{\sqrt{1+|y|^{2}}}\cdot\nabla\partial_{y_{k}}v_{\text{re}}-\frac{2+|y|^{2}}{(1+|y|^{2})^{3/2}}\partial_{y_{k}}v_{\text{re}}+\frac{y}{\sqrt{1+|y|^{2}}}\partial_{y_{k}}q.$
@@
The other terms on the right-hand side of (3.12) can be seen belong to $L^{2}(\mathbb{R}^{2})$. Applying Lemma 3.2, we know that
@@ 3.4. Pointwise Decay estimates for the self-similar solution
We have the following lemma, which plays an important role in improving decay estimates for weak solutions of equation (3.1). This kind of lemma already appeared in [18, Lemma 4.1] in the 3D case; see also [28, Lemma 3.15]. We present the lemma for the 2D case; the proof is essentially the same.
@@
(3.18) $|\nabla\Theta(x)|$ $\leq\int_{0}^{1}s^{-\frac{1}{2}}\int_{\mathbb{R}^{2}}\frac{1}{(\sqrt{1-s}+|x-y|)^{3}}\frac{1}{(\sqrt{s}+|y|)^{2}}dyds$
@@
$|\nabla\Theta(x)|\leq\frac{C}{1+|x|^{2}}$
@@
$|\nabla v_{\textrm{re}}(x)|$ $\leq\left|\int_{0}^{1}\int_{\mathbb{R}^{2}}\nabla S(x-y,1-s)s^{-\frac{3}{2}}f\left(\frac{y}{\sqrt{s}}\right)dyds\right|$
@@
If the initial data has higher regularity with $\|u_{0}\|_{C^{2}(S^{1})}<+\infty$, then we can apply [8, Lemma 4] to the vector field $F=v_{0}\otimes v_{0}=e^{\Delta}u_{0}\otimes e^{\Delta}u_{0}$, to conclude that
@@
at time $t=1$ is bounded by $C(|u_{0}\|_{C^{2}(S^{1})})(1+|x|)^{-3}$. The proof makes essential use of the cancellation properties of the Oseen kernel. On the other hand, all the other terms in $(v_{\textrm{re}}+v_{0})\otimes(v_{\textrm{re}}+v_{0})$ except $v_{0}\otimes v_{0}$ has bound $C(1+|x|)^{-3}$ at least. And their contribution in (3.19) can be bound by
@@ 4. Proof of the main theorem
Let divergence-free initial data $u_{0}$ on $\mathbb{R}^{2}$ be self-similar, and that $\int_{\partial B_{R}}u_{0}\cdot nd\sigma=0$,
@@
Assume $u(x,t)$ be a self-similar solution to (1.1) in the sense of Definition 1.1. Then $v(x)=u(x,1)$ the solution profile at time t = 1, belongs to $C^{\infty}(\mathbb{R}^{2})$. Let $v_{0}=e^{\Delta}u_{0}$ and $v_{\textrm{re}}(x)=v(x)-v_{0}(x)$, then we have
@@
We first set up the framework for the application of the Leray-Schauder fixed point theorem. Define the function space
(4.1) $X=\{{\boldsymbol{v}}:\ {\rm div\thinspace}{\boldsymbol{v}}=0,\ {\boldsymbol{v}}\in C_{loc}(\mathbb{R}^{2})\},$
equipped with the norm
$\|{\boldsymbol{v}}\|_{X}=\|\sqrt{1+|x|^{2}}{\boldsymbol{v}}(x)\|_{C(\mathbb{R}^{2})}.$
The function space $X$ is a Banach space. For each ${\boldsymbol{v}}\in X$, we define its self-similar extension by
@@
We now define an operator $K:X\times[0,1]\to X$ by solving the nonhomogeneous Stokes system and restrict the solution at time $t=1$:
@@
$|K({\boldsymbol{v}},\sigma)(x,1)|\leq\frac{C(A,M)}{1+|x|^{2}}\leq C(A,M)$
@@
Furthermore, $K$ is compact because its main term $\Phi[(\sigma e^{t\Delta}u_{0})\otimes(\sigma e^{t\Delta}u_{0})]\big|_{t=1}=\sigma^{2}\Phi[(e^{t\Delta}u_{0})\otimes(e^{t\Delta}u_{0})]\big|_{t=1}$ has a one-dimensional range. While the other terms of $K$ decay at infinity at least with the rate $\frac{1}{1+|x|^{2}}$ by Theorem 4.1. And they possess at least Hölder continuity in any bounded region due to classical elliptic estimates for the linear Stokes operator.
@@
    It is uniquely solvable in $X$ for small $\sigma$ by classical result; see for example [10] and [8, Proposition 3]
@@
By the Leray-Schauder theorem, there is a solution ${\boldsymbol{v}}\in X$ of (4.3) with $\sigma=1$. It follows that $u_{\textrm{re}}=E{\boldsymbol{v}}$ satisfies the nonhomogeneous Stokes system (3.13) with $f=-(e^{t\Delta}u_{0}+u_{\textrm{re}})\otimes(e^{t\Delta}u_{0}+u_{\textrm{re}})$, and hence $u=e^{t\Delta}u_{0}+u_{\textrm{re}}$ is a self-similar solution of (1.1) with initial data $u_{0}$. It is easy to check the solution we find satisfies the Definition 1.1. Next, we prove the pointwise estimates in Theorem 1.1. We have by Theorem 4.1 and Lemma 2.1 that
@@
In this appendix, we prove Lemma 2.1. Before we proceed, we give the following proposition concerning the difference estimates for (1)(-1)-homogeneous, locally Hölder continuous vector field, which is needed in the proof of Lemma 2.1.
@@
Let $u_{0}(x)$ be a (1)(-1)-homogeneous vector field in $\mathbb{R}^{n}\setminus\{0\}$ ($n\geq 2$) such that its restriction to the unit sphere satisfies $u_{0}\in C^{0,\beta}(S^{n-1})$ with $0<\beta<1$. Let $A=\|u_{0}\|_{C^{0,\beta}(S^{n-1})}$. Then for any $y,z\in\mathbb{R}^{n}\setminus\{0\}$, the following estimate holds:
(A.1) $|u_{0}(z)-u_{0}(y)|\leq C(A)\frac{|z-y|^{\beta}}{\min(|z|,|y|)^{1+\beta}}.$
@@
Let $y,z\in\mathbb{R}^{n}\setminus\{0\}$. We use polar coordinates $y=\rho\theta$ and $z=r\omega$, where $\rho=|y|$, $r=|z|$, and $\theta,\omega\in S^{n-1}$. By the (1)(-1)-homogeneity of $u_{0}$, we have $u_{0}(y)=\rho^{-1}u_{0}(\theta)$ and $u_{0}(z)=r^{-1}u_{0}(\omega)$. The difference is decomposed as:
$|u_{0}(z)-u_{0}(y)|$ $=\left|\frac{1}{r}u_{0}(\omega)-\frac{1}{\rho}u_{0}(\theta)\right|$
$=\left|\frac{1}{r}(u_{0}(\omega)-u_{0}(\theta))+\left(\frac{1}{r}-\frac{1}{\rho}\right)u_{0}(\theta)\right|$
$\leq\frac{1}{r}|u_{0}(\omega)-u_{0}(\theta)|+|u_{0}(\theta)|\left|\frac{1}{r}-\frac{1}{\rho}\right|=:\text{Term I}+\text{Term II}.$
Step 1: Estimating Term I (Angular Term). Since $u_{0}\in C^{\beta}(S^{n-1})$, we have $|u_{0}(\omega)-u_{0}(\theta)|\leq A|\omega-\theta|^{\beta}$. Utilizing the geometric inequality
$\left|\omega-\theta\right|=\left|\frac{z}{|z|}-\frac{y}{|y|}\right|\leq\frac{2|z-y|}{\max(|z|,|y|)},$
@@
$\frac{1}{r}|u_{0}(\omega)-u_{0}(\theta)|\leq\frac{A}{r}\left(\frac{2|z-y|}{\max(r,\rho)}\right)^{\beta}\leq\frac{C(A)}{\min(r,\rho)^{1+\beta}}|z-y|^{\beta}.$
Step 2: Estimating Term II (Radial Term). Without loss of generality, assume $\rho\leq r$. We prove Term II is bounded by $C\rho^{-(1+\beta)}|z-y|^{\beta}$ by considering two cases for $|r-\rho|$:
Case A: Near-field radial difference ($|r-\rho|\leq\frac{\rho}{2}$). In this case, $r\geq\rho$ and $|r-\rho|\leq|z-y|$. We have:
$\left|\frac{1}{r}-\frac{1}{\rho}\right|=\frac{|r-\rho|}{r\rho}=\frac{|r-\rho|^{\beta}\cdot|r-\rho|^{1-\beta}}{r\rho}\leq\frac{|z-y|^{\beta}\cdot(\rho/2)^{1-\beta}}{\rho^{2}}=\frac{1}{2^{1-\beta}}\frac{|z-y|^{\beta}}{\rho^{1+\beta}}.$
Case B: Far-field radial difference ($|r-\rho|>\frac{\rho}{2}$). In this case, $|z-y|\geq|r-\rho|>\frac{\rho}{2}$, which implies $1<\frac{2^{\beta}|z-y|^{\beta}}{\rho^{\beta}}$. Since $r\geq\rho$, we have:
$\left|\frac{1}{r}-\frac{1}{\rho}\right|\leq\frac{1}{r}+\frac{1}{\rho}\leq\frac{2}{\rho}<\frac{2}{\rho}\cdot\frac{2^{\beta}|z-y|^{\beta}}{\rho^{\beta}}=\frac{2^{1+\beta}|z-y|^{\beta}}{\rho^{1+\beta}}.$
Combining both cases, we have $\text{Term II}\leq\frac{CA}{\min(r,\rho)^{1+\beta}}|z-y|^{\beta}$. Summing the estimates for Term I and Term II yields (A.1). ∎
@@
$I_{2}$ $\leq\int_{\left\{z:|y-z|\leq 3|y|\right\}}e^{-\frac{|y-z|^{2}}{4}}\frac{C(A)}{|y|}dz$
$\leq\frac{C(A)}{|y|}\int_{\left\{|\tilde{z}|\leq 3|y|\right\}}e^{-\frac{|\tilde{z}|^{2}}{4}}dz\leq\frac{C(A)}{|y|}.$
@@
$I_{3}$ $\leq\frac{C(A)}{|y|}\int_{|z|\geq 2|y|}e^{-\frac{|y-z|^{2}}{8}}e^{-\frac{|y-z|^{2}}{8}}dz$
$\leq\frac{C(A)}{|y|}e^{-\frac{|y|^{2}}{8}}\int_{|z|\geq 2|y|}e^{-\frac{|z|^{2}}{32}}dz\leq\frac{C(A)}{|y|}e^{-\frac{|y|^{2}}{8}}.$
@@
The case $k=1$. Next, we prove (2.12) for $k=1$. We consider first the component $\beta=1$, and we have that
@@
$|\partial_{i}v_{0}(y)|$ $\leq\int_{\mathbb{R}^{2}}e^{-\frac{|y-z|^{2}}{4}}|y_{i}-z_{i}|\frac{A}{|z|}dz$
$\leq\int_{\mathbb{R}^{2}}e^{-\frac{|y-z|^{2}}{4}}|y-z|\frac{A}{|z|}dz.$
@@
Next, we consider the case for $k=1$ and $0<\beta<1$. The derivative $\partial_{i}v_{0}(y)$ can be expressed as:
@@
$\left|\int_{L_{2}}\partial_{y_{i}}G(y-z)[u_{0}(z)-u_{0}(y)]\,dz\right|$ $\leq\frac{C(A)}{|y|^{1+\beta}}\int_{|\tilde{z}|\leq 3|y|}|\partial_{i}G(\tilde{z})||\tilde{z}|^{\beta}\,d\tilde{z}$
$\leq\frac{C(A,\beta)}{|y|^{1+\beta}}.$
@@
$\left|\int_{L_{3}}\partial_{y_{i}}G(y-z)u_{0}(z)\,dz-u_{0}(y)\int_{L_{3}}\partial_{y_{i}}G(y-z)\,dz\right|\leq C(A)e^{-\frac{|y|^{2}}{16}}$
Combining the estimates for $L_{1}$, $L_{2}$ and $L_{3}$, we obtain that for $\beta\in(0,1)$:

### v2

@@ 3.3. Weighted $H^{2}$-Estimates of $\nabla v_{\textnormal{re}}$
$\left\{\begin{aligned} &-\Delta V+\overline{V}\cdot\nabla V-\dfrac{1}{2}(y\cdot\nabla V+V)+\nabla P=f,\\ &{\rm div\thinspace}V=0,\end{aligned}\right.\text{in }\mathbb{R}^{2}.$
@@
(3.11) $-\Delta v_{\text{re}}$ $=\frac{1}{2}y\cdot\nabla v_{\text{re}}+\frac{1}{2}v_{\text{re}}-\mathbb{P}(v_{\text{re}}\cdot\nabla v_{\text{re}}+v_{0}\cdot\nabla v_{\text{re}}+v_{\text{re}}\cdot\nabla v_{0}+v_{0}\cdot\nabla v_{0})$
@@
Let $V_{k}=\sqrt{1+|y|^{2}}\partial_{y_{k}}v_{\text{re}}$ and $q_{k}=\sqrt{1+|y|^{2}}\partial_{y_{k}}q$ for $k=1,2$, we can find that
@@
$+v_{\text{re}}\cdot\frac{y}{\sqrt{1+|y|^{2}}}\partial_{y_{k}}v_{\text{re}}-\frac{1}{2}V_{k}-\frac{|y|^{2}}{2\sqrt{1+|y|^{2}}}\partial_{y_{k}}v_{\text{re}}$
$-2\frac{y}{\sqrt{1+|y|^{2}}}\cdot\nabla\partial_{y_{k}}v_{\text{re}}-\frac{2+|y|^{2}}{(1+|y|^{2})^{3/2}}\partial_{y_{k}}v_{\text{re}}+\frac{y}{\sqrt{1+|y|^{2}}}\partial_{y_{k}}q.$
@@
The other terms on the right-hand side of (3.12) can be seen belong to $L^{2}(\mathbb{R}^{2})$. Applying Lemma 3.2 to (3.12), we know that
@@ 3.4. Pointwise Decay estimates for the self-similar solution
We have the following lemma, which plays an important role in improving decay estimates for weak solutions of equation (3.1). This kind of lemma already appeared in [25, Lemma 4.1] in the 3D case; see also [48, Lemma 3.15]. We present the lemma for the 2D case; the proof is essentially the same.
@@
(3.18) $|\nabla\Theta(x)|$ $\leq\int_{0}^{1}s^{-\frac{1}{2}}\int_{\mathbb{R}^{2}}\frac{1}{(\sqrt{1-s}+|x-y|)^{3}}\frac{1}{(\sqrt{s}+|y|)^{2}}dyds.$
@@
$|\nabla\Theta(x)|\leq\frac{C}{1+|x|^{2}},$
@@
(3.20) $|\nabla v_{\textrm{re}}(x)|$ $\leq\left|\int_{0}^{1}\int_{\mathbb{R}^{2}}\nabla S(x-y,1-s)s^{-\frac{3}{2}}f\left(\frac{y}{\sqrt{s}}\right)dyds\right|$
@@
If the initial data has higher regularity with $\|u_{0}\|_{C^{2}(S^{1})}<+\infty$, then we can apply [12, Lemma 4] to the vector field $F=v_{0}\otimes v_{0}=e^{\Delta}u_{0}\otimes e^{\Delta}u_{0}$, to conclude that
@@
at time $t=1$ is bounded by $C(|u_{0}\|_{C^{2}(S^{1})})(1+|x|)^{-3}\ln(1+|x|)$. The proof makes essential use of the cancellation properties of the Oseen kernel. On the other hand, all the other terms in $(v_{\textrm{re}}+v_{0})\otimes(v_{\textrm{re}}+v_{0})$ except $v_{0}\otimes v_{0}$ has bound $C(1+|x|)^{-3}$ at least. And their contribution in (3.19) can be calculated similarly to that of Lemma 3.3 and bound by
@@ 4. Proof of the main theorem
Let divergence-free initial data $u_{0}$ on $\mathbb{R}^{2}$ be self-similar, and that $\int_{\partial S^{1}}u_{0}\cdot nd\sigma=0$,
@@
Assume $u(x,t)$ is a self-similar solution to (1.1) in the sense of Definition 1.1. Then $v(x)=u(x,1)$ the solution profile at time t = 1, belongs to $C^{\infty}(\mathbb{R}^{2})$. Let $v_{0}=e^{\Delta}u_{0}$ and $v_{\textrm{re}}(x)=v(x)-v_{0}(x)$, assume $v_{\textrm{re}}\in X$, then we have
@@
We also recall that
(4.1) $X=\{{\boldsymbol{v}}:\ {\rm div\thinspace}{\boldsymbol{v}}=0,\ {\boldsymbol{v}}\in C^{1}(\mathbb{R}^{2}),~\|{\boldsymbol{v}}\|_{X}<+\infty\},$
and the norm $\|\cdot\|_{X}$ is
$\|{\boldsymbol{v}}\|_{X}=\|(1+|x|^{2}){\boldsymbol{v}}\|_{C(\mathbb{R}^{2})}+\|(1+|x|^{2})\nabla{\boldsymbol{v}}\|_{C(\mathbb{R}^{2})}.$
@@
We try apply the Leray-Schauder fixed point theorem. For each ${\boldsymbol{v}}\in X$, we define its self-similar extension by
@@
We now define an operator $K:X\times[0,1]\to X$ by solving the nonhomogeneous Stokes system and restricting the solution at time $t=1$:
@@
$|K({\boldsymbol{v}},\sigma)(x,1)|,~|\nabla K({\boldsymbol{v}},\sigma)(x,1)|\leq\frac{C(A,M)}{1+|x|^{2}}\leq C(A,M)$
@@
Furthermore, $K$ is compact because its main term $\Phi[(\sigma e^{t\Delta}u_{0})\otimes(\sigma e^{t\Delta}u_{0})]\big|_{t=1}=\sigma^{2}\Phi[(e^{t\Delta}u_{0})\otimes(e^{t\Delta}u_{0})]\big|_{t=1}$ has a one-dimensional range. While the other terms of $(\sigma e^{t\Delta}u_{0}+E{\boldsymbol{v}})\otimes(\sigma e^{t\Delta}u_{0}+E{\boldsymbol{v}})$ except $\sigma e^{t\Delta}u_{0}\otimes\sigma e^{t\Delta}u_{0}$ decay with rate faster or equal to $\frac{1}{1+|x|^{3}}$. Similar to calculations in Lemma 3.3, we know that $\Phi[\sigma e^{t\Delta}u_{0}\otimes E{\boldsymbol{v}}+E{\boldsymbol{v}}\otimes(\sigma e^{t\Delta}u_{0}+E{\boldsymbol{v}})]\big|_{t=1}$ decay at infinity at least with the rate $\frac{1}{1+|x|^{3}}$. Due to the same reason, we see that $\nabla\Phi[\sigma e^{t\Delta}u_{0}\otimes E{\boldsymbol{v}}+E{\boldsymbol{v}}\otimes(\sigma e^{t\Delta}u_{0}+E{\boldsymbol{v}})]\big|_{t=1}$ decay at infinity at least with the rate $\frac{1}{1+|x|^{3}}$. And they possess at least $C^{1,\alpha}$ ($\alpha>0$) regularity in any bounded region due to classical elliptic estimates for the linear Stokes operator. It is then not hard to see that $K$ is compact. We can also easily verify that $K$ is continuous based on similar estimates to Lemma 3.3.
@@
    It is uniquely solvable in $X$ for small $\sigma$ by classical result; see for example [14] and [12, Proposition 3].
@@
By the Leray-Schauder theorem, there is a solution $v_{\textrm{re}}\in X$ of (4.3) with $\sigma=1$. It follows that $u_{\textrm{re}}=Ev_{\textrm{re}}$ satisfies the nonhomogeneous Stokes system (3.13) with $f=-(e^{t\Delta}u_{0}+u_{\textrm{re}})\otimes(e^{t\Delta}u_{0}+u_{\textrm{re}})$, and hence $u=e^{t\Delta}u_{0}+u_{\textrm{re}}$ is a self-similar solution of (1.1) with initial data $u_{0}$. It is easy to check the solution we find satisfies the Definition 1.1. Next, we prove the pointwise estimates in Theorem 1.1. We have by Theorem 3.1 that
@@
In this appendix, we prove Lemma 2.1. Before we proceed, we give the following proposition concerning the estimates for (1)(-1)-homogeneous, locally Hölder continuous vector field in any dimension, which is needed in the proof of Lemma 2.1.
@@
Let $u_{0}(x)$ be a (1)(-1)-homogeneous vector field in $\mathbb{R}^{n}\setminus\{0\}$ ($n\geq 2$) such that its restriction to the unit sphere satisfies $u_{0}\in C^{0,\beta}(S^{n-1})$ with $0<\beta<1$. Let $A=\|u_{0}\|_{C^{0,\beta}(S^{n-1})}$. Then for any $y,\tilde{y}\in\mathbb{R}^{n}\setminus\{0\}$, the following estimate holds:
(A.1) $|u_{0}(y)-u_{0}(\tilde{y})|\leq C(A)\frac{|y-\tilde{y}|^{\beta}}{\min(|y|,|\tilde{y}|)^{1+\beta}}.$
@@
[…]
$=\left|\frac{1}{\tilde{\rho}}(u_{0}(\theta)-u_{0}(\tilde{\theta}))+\left(\frac{1}{\rho}-\frac{1}{\tilde{\rho}}\right)u_{0}(\theta)\right|$
$\leq\frac{1}{\tilde{\rho}}|u_{0}(\theta)-u_{0}(\tilde{\theta})|+|u_{0}(\theta)|\left|\frac{1}{\rho}-\frac{1}{\tilde{\rho}}\right|=:\textrm{I}+\textrm{II}.$
Step 1: Estimating Term I. Since $u_{0}\in C^{\beta}(S^{n-1})$, we have $u_{0}(\theta)-|u_{0}(\tilde{\theta})|\leq A|\theta-\tilde{\theta}|^{\beta}$. Utilizing the geometric inequality
$\left|\theta-\tilde{\theta}\right|=\left|\frac{y}{|y|}-\frac{\tilde{y}}{|\tilde{y}|}\right|\leq\frac{2|y-\tilde{y}|}{\max(|y|,|\tilde{y}|)}$,
@@
$\frac{1}{\tilde{\rho}}|u_{0}(\tilde{\theta})-u_{0}(\theta)|\leq\frac{A}{\tilde{\rho}}\left(\frac{2|y-\tilde{y}|}{\max(\tilde{\rho},\rho)}\right)^{\beta}\leq\frac{C(A)}{\min(\tilde{\rho},\rho)^{1+\beta}}|y-\tilde{y}|^{\beta}.$
[…]
$\left|\frac{1}{\tilde{\rho}}-\frac{1}{\rho}\right|=\frac{|\rho-\tilde{\rho}|}{\tilde{\rho}\rho}=\frac{|\rho-\tilde{\rho}|^{\beta}\cdot|\rho-\tilde{\rho}|^{1-\beta}}{\tilde{\rho}\rho}\leq\frac{|y-\tilde{y}|^{\beta}\cdot(\rho/2)^{1-\beta}}{\rho^{2}}=\frac{1}{2^{1-\beta}}\frac{|y-\tilde{y}|^{\beta}}{\rho^{1+\beta}}.$
Case B: Far-field radial difference ($|\tilde{\rho}-\rho|>\frac{\rho}{2}$). In this case, $|y-\tilde{y}|\geq|\tilde{\rho}-\rho|>\frac{\rho}{2}$, which implies $1<\frac{2^{\beta}|y-\tilde{y}|^{\beta}}{\rho^{\beta}}$. Since $\tilde{\rho}\geq\rho$, we have
$\left|\frac{1}{\tilde{\rho}}-\frac{1}{\rho}\right|\leq\frac{1}{\tilde{\rho}}+\frac{1}{\rho}\leq\frac{2}{\rho}<\frac{2}{\rho}\cdot\frac{2^{\beta}|y-\tilde{y}|^{\beta}}{\rho^{\beta}}=\frac{2^{1+\beta}|y-\tilde{y}|^{\beta}}{\rho^{1+\beta}}.$
Combining both cases, we have  $\textrm{ II}\leq\frac{CA}{\min(\tilde{\rho},\rho)^{1+\beta}}|y-\tilde{y}|^{\beta}$. Summing the estimates for terms I and II yields (A.1). ∎
@@
$I_{2}$ $\leq\int_{\left\{z:|y-z|\leq 3|y|\right\}}e^{-\frac{|y-z|^{2}}{4}}\frac{C(A)}{|y|}dz\leq\frac{C(A)}{|y|}\int_{\left\{|\tilde{z}|\leq 3|y|\right\}}e^{-\frac{|\tilde{z}|^{2}}{4}}dz\leq\frac{C(A)}{|y|}.$
@@
$I_{3}$ $\leq\frac{C(A)}{|y|}\int_{|z|\geq 2|y|}e^{-\frac{|y-z|^{2}}{8}}e^{-\frac{|y-z|^{2}}{8}}dz\leq\frac{C(A)}{|y|}e^{-\frac{|y|^{2}}{8}}\int_{|z|\geq 2|y|}e^{-\frac{|z|^{2}}{32}}dz\leq\frac{C(A)}{|y|}e^{-\frac{|y|^{2}}{8}}.$
@@
The case $k=1$. Next, we prove (2.15) for $k=1$. The case for $\beta=0$ is trivial. We then consider the case when the component $\beta=1$. We have that
@@
$|\partial_{i}v_{0}(y)|$ $\leq\int_{\mathbb{R}^{2}}e^{-\frac{|y-z|^{2}}{4}}|y_{i}-z_{i}|\frac{A}{|z|}dz\leq\int_{\mathbb{R}^{2}}e^{-\frac{|y-z|^{2}}{4}}|y-z|\frac{A}{|z|}dz.$
@@
Next, we consider the case for $k=1$ and $0<\beta<1$. The derivative $\partial_{i}v_{0}(y)$ can be expressed as
@@
$\left|\int_{L_{2}}\partial_{y_{i}}G(y-z)[u_{0}(z)-u_{0}(y)]\,dz\right|$ $\leq\frac{C(A)}{|y|^{1+\beta}}\int_{|\tilde{z}|\leq 3|y|}|\partial_{i}G(\tilde{z})||\tilde{z}|^{\beta}\,d\tilde{z}\leq\frac{C(A,\beta)}{|y|^{1+\beta}}.$
@@
$\left|\int_{L_{3}}\partial_{y_{i}}G(y-z)u_{0}(z)\,dz-u_{0}(y)\int_{L_{3}}\partial_{y_{i}}G(y-z)\,dz\right|\leq C(A)e^{-\frac{|y|^{2}}{16}}.$
Combining the estimates for $L_{1}$, $L_{2}$ and $L_{3}$, we obtain that for $\beta\in(0,1)$,